OpenStudy (anonymous):

Suppose the slope of line AB is greater than 0 and less than 1. What inequality represents the slope of a line perpendicular to line AB? (Hint: This part of the question has one right answer.) Give an example of the slopes for a pair of lines that would satisfy both inequalities. (Many examples are possible.) This is a topic on a class discussion, and I honestly don't understand it. I need help mainly on the second part of writing an example rather than the first part. Please help=)

OpenStudy (anonymous):

\[0<m_1<1\] when lines are perpendicular, their slopes are negative reciprocals so, \[ m_2=-{1\over m_1}\\ 0<m_1<1\\ \infty>{1\over m_1}>1\\ \boxed{-\infty<m_2<-1} \]

OpenStudy (anonymous):

So that is one possible answer for the second half?

OpenStudy (anonymous):

@Nurali the solution you've posted seems controvercial

OpenStudy (anonymous):

so, a line perpendicular to AB with slope of \(m_1\), has a slope of \(m_2\). then we follow the above steps.

OpenStudy (anonymous):

it is given that \(0<m_1<1\)

OpenStudy (anonymous):

so, we flip everything.. i.e., take the reciprocal. and the inequality signs flip \[{1\over0}>{1\over m_1}>{1\over 1}\\ \implies \infty>{1\over m_1}>1\]

OpenStudy (anonymous):

if we multiply the inequality by -1, the signs flip again \[-\infty<{-1\over m_1}<-1\]

OpenStudy (anonymous):

but, by the property of perpendicular lines, their slopes are negative reciprocals of eachother. i.e., \(m_2=\Large {-1\over m_1}\)

OpenStudy (anonymous):

so, we finally get: \[-\infty <m_2<-1\]

OpenStudy (anonymous):

kapeesh?

OpenStudy (anonymous):

I wish, I feel rather stupid right now haha. I honestly don't understand. I know this is a correct answer, it is an example my teacher posted but I don't understand how to create one myself y = 1/2 x + 10 and y = -2 x + 15

OpenStudy (anonymous):

its ok. did you follow the steps I posted up here one by one? the idea behind this is tis: \(\mathbf{\text{if two lines with slopes m_1 and m_2 are peprendicular,}}\\ \mathbf{\text{ then m_1= -1/m_2}}\)

OpenStudy (anonymous):